# MSTN: Fast and Efficient Multivariate Time Series Model

### v1

@@ 4.1.1 Architecture Overview
The proposed MSTN (Fig. 1a) is a novel hybrid DL architecture founded on a hierarchical multi-scale modeling principle. It uses a dual-branch processing pipeline: (i) a convolutional branch that extracts detailed local motion patterns, and (ii) sequence modeling core that learns temporal dependencies in the time-series data, which we empirically validate using both BiLSTM and Transformer variant in forecasting, imputation, classification and generalizability studies. (iii) The outputs of both branches are integrated using a learnable gated fusion mechanism, followed by channel-wise recalibration through SE blocks and MHTA, enabling the model to focus on critical, risk-indicating time steps. The fused representation is normalized, regularized with dropout, and linearly projected to produce probabilities. Training takes advantage of focal loss to mitigate class imbalance. Fig. 1b shows the processing pipeline of time series signals. CNN pathway captures local patterns, while BilSTM and Transformer captures long-range temporal dependencies in the time-series data, with gated fusion adaptively combining both feature streams.
4.1.2 Model Structure
As aforementioned, based on the multi-scale nature of time series, we propose MSTN with a parallel architecture to capture temporal patterns at different resolutions. For comprehensive temporal modeling, we design dual pathways within MSTN that process the time series through complementary mechanisms and simultaneously model both local and global variations through a parameter-efficient novel hybrid block. The MSTN architecture and signal processing pipeline are illustrated in Fig. 1.
@@ 4.1.2 Model Structure
We consider the following problem: Given multivariate time series samples $X_{1:L}\in\mathbb{R}^{L\times M}$ with lookback window sequence length $L$ and $M$ variates features, MSTN learns rich temporal representations for multiple tasks including forecasting, imputation and classification through a novel parallel architecture.
Forward Process. The input $X\in\mathbb{R}^{B\times T\times D}$ is processed through dual complementary pathways that operate simultaneously. The CNN pathway captures local temporal patterns while the Transformer pathway models global dependencies or BiLSTM models sequential intelligence, with their outputs fused through sophisticated gated mechanisms for enhanced representation learning. To capture multi-scale temporal variations in length-$T$ time series with $D$ variates organized as $X_{1D}\in\mathbb{R}^{B\times T\times D}$, the architecture processes inputs through two parallel pathways with convolutional pathway and sequence modeling (Transformer or BiLSTM pathway).
The convolutional pathway captures fine-grained, short-term temporal patterns through hierarchical feature learning. The input undergoes successive 1D convolutions to extract multi-scale features, expanding the feature dimension while preserving temporal resolution, followed by refinement of the representations. To aggregate these temporal features into a compact representation, we apply global average pooling across the entire sequence:
$H_{\text{conv}}^{(1)}$ $=\text{ReLU}(\text{Conv1D}_{7}(X_{1D}))\in\mathbb{R}^{B\times 128\times T}$ (1)
$H_{\text{conv}}^{(2)}$ $=\text{ReLU}(\text{Conv1D}_{5}(H_{\text{conv}}^{(1)}))\in\mathbb{R}^{B\times 64\times T}$ (2)
@@
where each of the 64 feature channels is summarized by its average activation over time through temporal pooling. The $\text{Conv1D}_{k}$ convolutional layers employ a hierarchical design where the first layer with kernel size $k=7$ extracts basic temporal patterns, and the second layer with kernel size $k=5$ learns combinations of these patterns into more complex features, with padding to maintain temporal dimensions, and batch normalization stabilizes training.
We empirically validate two complementary sequence modeling approaches: MSTN-Transformer and MSTN-BiLSTM. Features from both pathways are elegantly concatenated, uniting the CNN’s local precision with either Transformer or BiLSTM, to capture long-range temporal dependencies. When using Transformer: The architecture captures long-range dependencies through a multi-layer transformer encoder. The input is processed through self-attention mechanisms where each of the T time steps maintains a 64-dimensional representation enriched with global contextual information. To consolidate these temporal representations, we apply sequence mean pooling across the time dimension:
$H_{\text{trans}}=\text{TransformerEnc}(X)\in\mathbb{R}^{B\times T\times 64};\quad\mathbf{z}_{\text{trans}}=\frac{1}{T}\sum_{t=1}^{T}H_{\text{trans},t}\in\mathbb{R}^{B\times 64}$ (4)
The transformer encoder uses 4 layers with 8 attention heads. When using BiLSTM: The architecture captures bidirectional sequential dependencies, processing sequences in both forward and backward directions to learn complex temporal dynamics and memorize long-range patterns:
$H_{\text{bilstm}}=\text{BiLSTM}(X)\in\mathbb{R}^{B\times T\times 128};\quad\mathbf{z}_{\text{bilstm}}=\frac{1}{T}\sum_{t=1}^{T}H_{\text{bilstm},t}\in\mathbb{R}^{B\times 128}$ (5)
The BiLSTM employs 2 layers with 64 hidden units per direction, capturing both causal and anti-causal temporal relationships. The CNN pathway utilizes global average pooling to capture local pattern summaries, while the Transformer and BiLSTM pathways employ sequence mean pooling to preserve global contextual information. The parallel features undergo sophisticated fusion through a multi-stage enhancement process that intelligently combines the strengths of both pathways. For MSTN-Transformer:
$\mathbf{z}_{\text{concat}}=[\mathbf{z}_{\text{cnn}};\mathbf{z}_{\text{trans}}]\in\mathbb{R}^{128}$ (6)
where $\mathbf{z}_{\text{cnn}}\in\mathbb{R}^{64}$ and $\mathbf{z}_{\text{trans}}\in\mathbb{R}^{64}$. For MSTN-BiLSTM:
$\mathbf{z}_{\text{concat}}=[\mathbf{z}_{\text{cnn}};\mathbf{z}_{\text{bilstm}}]\in\mathbb{R}^{192}$ (7)
where $\mathbf{z}_{\text{cnn}}\in\mathbb{R}^{64}$ and $\mathbf{z}_{\text{bilstm}}\in\mathbb{R}^{128}$. The architecture then performs adaptive weighting through an intelligent gated fusion mechanism, illustrated in Fig. 1, that learns to dynamically balance feature contributions:
@@
where for MSTN-Transformer: $W_{g}\in\mathbb{R}^{128\times 128}$, $b_{g}\in\mathbb{R}^{128}$ and for MSTN-BiLSTM: $W_{g}\in\mathbb{R}^{192\times 192}$, $b_{g}\in\mathbb{R}^{192}$. This mechanism learns to emphasize the most relevant features from each pathway, creating a harmonious blend of multi-scale temporal intelligence. The fused features are reshaped for sequence processing: $z_{\text{fused\_seq}}=z_{\text{fused}}\text{.unsqueeze}(1)\text{.repeat}(1,T,1)\in\mathbb{R}^{B\times T\times d}$ where $d=128$ for MSTN-Transformer and $d=192$ for MSTN-BiLSTM. Channel-wise attention then enhances features through separate SE blocks for each variant: For MSTN-Transformer: $z_{\text{se}}=z_{\text{fused\_seq}}\odot\sigma(W_{2}\text{ReLU}(W_{1}\frac{1}{T}\sum_{t=1}^{T}z_{\text{fused\_seq},t}))$ where $W_{1}\in\mathbb{R}^{16\times 128}$, $W_{2}\in\mathbb{R}^{128\times 16}$ with reduction ratio 8. For MSTN-BiLSTM: $z_{\text{se}}=z_{\text{fused\_seq}}\odot\sigma(W_{2}\text{ReLU}(W_{1}\frac{1}{T}\sum_{t=1}^{T}z_{\text{fused\_seq},t}))$ where $W_{1}\in\mathbb{R}^{24\times 192}$, $W_{2}\in\mathbb{R}^{192\times 24}$ with reduction ratio 8. The SE mechanism amplifies informative channels while suppressing less useful ones. Temporal relationships are refined through multi-head attention that captures interdependencies across the enhanced features:
@@
$\mathbf{z}_{\text{final}}=\text{Dropout}\left(\text{LayerNorm}\left(T^{-1}\textstyle\sum_{t=1}^{T}\text{MHA}(\mathbf{z}_{\text{se}})_{t}\right),p=0.3\right)\in\mathbb{R}^{B\times 192}$
with 4 attention heads (dim=64). The attention output undergoes layer normalization for stabilization and dropout for regularization, producing refined features $\mathbf{z}_{\text{final}}$ for the final classification head. The parallel pathway design enables MSTN to achieve comprehensive temporal modeling where $\mathcal{F}_{\text{CNN}}$ captures local patterns and $\mathcal{F}_{\text{Transformer}}$ models global dependencies, with hierarchical fusion creating synergistic representations that surpass individual pathway capabilities. This architectural innovation establishes a new paradigm for multi-scale time series analysis.
4.1.3 Method-Level Innovation
Many recent time-series models including PatchTST [nie2023timeseriesworth64] (patch-based tokenization), TimesNet [wu2023timesnettemporal2dvariationmodeling] (fixed frequency transformations), and LLM-based approaches like LLM4TS [10.1145/3719207] and TIME-LLM [jin2024timellmtimeseriesforecasting] (frozen backbone architectures) often rely on rigid, fixed-scale structural priors. This architectural inflexibility forces a fundamental trade-off: models either compromise on capturing multi-scale temporal diversity or suffer from high inference latency, rendering them suboptimal for real-time, resource-constrained deployment.
The proposed MSTN addresses this core challenge through a novel multi-scale hierarchical architecture designed to capture both local patterns and long-range temporal dependencies without structural compromises. The architecture integrates three key components: (i) a multi-scale convolutional encoder that constructs a hierarchical feature pyramid; (ii) a sequence temporal modeling core, which we empirically validate with both BiLSTM and Transformer variants, establishing a flexible foundation for future architectural advancements; and (iii) a gated fusion mechanism augmented with SE and MHTA for dynamic, context-aware feature integration. This hierarchical design enables seamless adaptation to diverse temporal modeling paradigms while maintaining architectural consistency. Our comprehensive experiments demonstrate that these design choices allow MSTN to achieve improved performance compared to SOTA baselines while supporting low-latency inference critical for edge deployment.
4.2 Dataset and Baselines

### v2

@@ 4 Methods
This section details the MSTN architecture and its methodological innovations, followed by the baseline models, and training configurations.
@@ 4.1.1 Architecture Overview
The proposed Multi-Scale Temporal Network (MSTN; Fig. 1a) is a hybrid deep-learning architecture designed around a parallel multi-scale modeling paradigm. MSTN processes temporal signals through two coordinated encoding pathways, each specialized for complementary aspects of temporal structure. A key design element is the use of an early temporal aggregation (ETA) mechanism, applied directly to the outputs of the dual encoders, which collapses the full input sequence of length $L$ into a single fixed-dimensional representation ($L\rightarrow 1$). This operation ensures that all subsequent refinement stages operate in constant time with respect to sequence length, thereby enabling efficient processing of long temporal inputs.
The network comprises three principal modules. (i) A multi-scale convolutional branch captures localized temporal patterns using hierarchically stacked convolutions; its output is aggregated via global average pooling, promoting robustness and translation invariance. (ii) A sequence modeling branch captures long-range dependencies, instantiated using either a bidirectional LSTM (MSTN-BiLSTM) or a Transformer encoder (MSTN-Transformer). To maintain computational efficiency, this branch is compressed through sequence mean pooling, yielding a global summary vector. (iii) A multi-scale fusion and refinement module integrates the pooled representations from the two branches. The features are concatenated and passed through a scale-gated fusion (SGF) mechanism, which adaptively modulates the contributions of each pathway. The fused representation is subsequently refined using channel-wise recalibration via squeeze–excitation blocks and multi-head attention, enabling the network to emphasize the most informative feature dimensions. The final representation is normalized, regularized through dropout, and projected to the task-specific output space. To address class imbalance, training is performed using a focal loss objective. Figure 1b provides an overview of the end-to-end processing pipeline, from parallel feature extraction to multi-scale fusion and refinement.
4.1.2 Methodology
@@
We consider the following problem: Given multivariate time series samples $X_{1:L}\in\mathbb{R}^{L\times C}$ with a lookback window sequence length $L$ and $C$ variate features, MSTN learns temporal representations through a novel parallel architecture. We detail the complete process below.
Forward Process. The input $X\in\mathbb{R}^{B\times L\times C}$ is processed through dual complementary pathways that operate simultaneously. The CNN pathway captures local temporal patterns, while the sequence modeling pathway (implemented as either a Transformer or a BiLSTM) captures long-range temporal dependencies. Their outputs are fused through a sophisticated gating and attention mechanism for enhanced representation learning. The input undergoes successive 1D convolutions to extract multi-scale features, adjusting the feature dimension while preserving temporal resolution, followed by refinement of the representations. The $\text{Conv1D}_{k}$ convolutional layers employ a hierarchical design where the first layer with kernel size $k=7$ extracts basic temporal patterns, and the second layer with kernel size $k=5$ learns combinations of these patterns into more complex features, with padding to maintain temporal dimensions, and batch normalization stabilizes training. To aggregate these temporal features into a compact representation, we apply global average pooling across the entire sequence:
$H_{\text{conv}}^{(1)}$ $=\text{ReLU}(\text{BN}(\text{Conv1D}_{7}(X^{T})))\in\mathbb{R}^{B\times 128\times L}$ (1)
$H_{\text{conv}}^{(2)}$ $=\text{ReLU}(\text{Conv1D}_{5}(H_{\text{conv}}^{(1)}))\in\mathbb{R}^{B\times 64\times L}$ (2)
@@
where each of the 64 feature channels is summarized by its average activation over time through temporal pooling.
We employ two sequence modeling approaches: MSTN-Transformer and MSTN-BiLSTM. In MSTN-Transformer, the architecture captures long-range dependencies through a single Transformer encoder block composed of 4 layers with 8 attention heads. The input is processed through self-attention mechanisms where each of the L time steps maintains a 128-dimensional representation enriched with global contextual information. To consolidate these temporal representations, we apply sequence mean pooling across the time dimension:
$H_{\text{trans}}=\text{TransformerEnc}(X)\in\mathbb{R}^{B\times L\times 128};\quad\mathbf{z}_{\text{trans}}=\frac{1}{L}\sum_{t=1}^{L}H_{\text{trans},t}\in\mathbb{R}^{B\times 128}$ (4)
In MSTN-BiLSTM: The architecture captures bidirectional sequential dependencies, processing sequences in both forward and backward directions to learn complex temporal dynamics and memorize long-range patterns:
$H_{\text{bilstm}}=\text{BiLSTM}(X)\in\mathbb{R}^{B\times L\times 128};\quad\mathbf{z}_{\text{bilstm}}=\frac{1}{L}\sum_{t=1}^{L}H_{\text{bilstm},t}\in\mathbb{R}^{B\times 128}$ (5)
The BiLSTM employs 2 layers with 64 hidden units per direction, capturing both causal and anti-causal temporal relationships.
The CNN pathway utilizes global average pooling, and the sequence modeling pathways employ sequence mean pooling. These pooling operations, along with the subsequent reshaping, collectively form the ETA module, which collapses the temporal dimension $L$ to $1$. This is the critical step that ensures subsequent operations happen at an $\mathcal{O}(1)$ inference cost. The parallel features undergo sophisticated fusion through a multi-stage enhancement process that combines the strengths of both pathways. For MSTN-Transformer:
$\mathbf{z}_{\text{concat}}=[\mathbf{z}_{\text{cnn}};\mathbf{z}_{\text{trans/bilstm}}]\in\mathbb{R}^{192}$ (6)
where $\mathbf{z}_{\text{cnn}}\in\mathbb{R}^{64}$ and $\mathbf{z}_{\text{trans}}\in\mathbb{R}^{128}$. The architecture then performs adaptive weighting through an SGF mechanism, illustrated in Fig. 1, that learns to dynamically balance feature contributions:
@@
where for both variants: $W_{g}\in\mathbb{R}^{192\times 192}$, $b_{g}\in\mathbb{R}^{192}$. This mechanism learns to emphasize the most relevant features from each pathway, creating a harmonious blend of multi-scale temporal features. The fused features are reshaped for efficient single-token processing: $\mathbf{z}_{\text{seq}}=\mathbf{z}_{\text{fused}}\otimes\mathbf{1}\in\mathbb{R}^{B\times 1\times 192}$. Channel-wise attention then enhances features through an SE block. Given the temporal dimension is already $L=1$, the global pooling step within SE is trivially executed: $\mathbf{z}_{\text{se}}=\mathbf{z}_{\text{seq}}\odot\sigma(W_{2}\text{ReLU}(W_{1}\mathbf{z}_{\text{seq}}))$ where $W_{1}\in\mathbb{R}^{24\times 192}$, $W_{2}\in\mathbb{R}^{192\times 24}$ with reduction ratio 8. The SE mechanism amplifies informative channels while suppressing less useful ones. Feature dependencies are refined through MHA. Since the temporal dimension is condensed ($L=1$), the MHA block functions here as a sophisticated global feature recalibration layer rather than a temporal mixer, confirming $\mathcal{O}(1)$ complexity:
@@
$\mathbf{z}_{\text{final}}=\text{Dropout}\left(\text{LayerNorm}\left(\sum_{t=1}^{1}\text{MHA}(\mathbf{z}_{\text{se}})_{t}\right),p=0.3\right)\in\mathbb{R}^{B\times 192}$ (8)
where 192 is the feature dimension for both MSTN-Transformer and MSTN-BiLSTM, with 4 attention heads (dim=48 per head). The refined features $\mathbf{z}_{\text{final}}$ are then passed to the final task-specific head (imputation, forecasting, or classification). The refined features $\mathbf{z}_{\text{final}}$ are then passed to the final task-specific head (imputation, forecasting, or classification). The prediction head consists of a single linear layer that maps the $192$-dimensional aggregated feature vector to the task-specific output shape.
Task-Specific Prediction and Loss Functions: We detail the prediction head ($\hat{Y}$ or $\hat{X}$) and the corresponding loss function $\mathcal{L}$ for each of the three tasks:
Time Series Imputation: The head reconstructs the input sequence $\hat{X}_{1:L}\in\mathbb{R}^{B\times L\times C}$, and the loss is calculated as masked MSE only over the observed values ($\Omega$ is the mask matrix):
$\hat{X}_{1:L}$ $=W_{i}\mathbf{z}_{\text{final}}+b_{i}\in\mathbb{R}^{B\times L\times C}$ (9)
$\mathcal{L}_{\text{imputation}}$ $=\frac{1}{|\Omega|}\sum_{i=1}^{B}\sum_{t=1}^{L}\sum_{c=1}^{C}\Omega_{i,t,c}\cdot\left(\hat{X}_{i,t,c}-X_{i,t,c}\right)^{2}$ (10)
Long-Term Forecasting: The head produces a forecast $\hat{Y}_{1:T}\in\mathbb{R}^{B\times T\times C}$ for a horizon $T$, and the objective is to minimize the MSE:
$\hat{Y}_{1:T}$ $=W_{f}\mathbf{z}_{\text{final}}+b_{f}\in\mathbb{R}^{B\times T\times C}$ (11)
$\mathcal{L}_{\text{forecast}}$ $=\frac{1}{2TC}\sum_{i=1}^{B}\sum_{t=1}^{T}\sum_{c=1}^{C}\left(\hat{Y}_{i,t,c}-Y_{i,t,c}\right)^{2}$ (12)
Time Series Classification: The head computes class probabilities $P\in\mathbb{R}^{B\times K}$ over $K$ classes, and the loss is the canonical Cross-Entropy (CE) Loss:
$P$ $=\text{Softmax}(W_{c}\mathbf{z}_{\text{final}}+b_{c})\in\mathbb{R}^{B\times K}$ (13)
$\mathcal{L}_{\text{classify}}$ $=-\frac{1}{B}\sum_{i=1}^{B}\sum_{k=1}^{K}Y_{i,k}\log(P_{i,k})$ (14)
where $Y$ is the one-hot encoded true label. This structure allows the MSTN framework to be universally applied to various multivariate time series tasks.
4.2 Method Level Innovation
Table 6: Analysis of model architectures by design characteristics. This table outlines how different state-of-the-art models align with the established Channel Interaction (CI, CD, CP) strategies based on the temporal and channel characteristics they model. MSTN is positioned as a hybrid design (CI+CD). (Asym.: Asymmetry, Lag.: Lagginess, Pol.: Polarity, Gw.: group-wise, Dyn.: Dynamism, Ms.: Multi-scale)
Strategy Mechanism Characteristic Method
Asym. Lag. Pol. Gw. Dyn. Ms.
CI - - - - - - - PatchTST
CI - - - - - - - CycleNet
CI - - - - - - - DLinear
CI - - - - - - - Timer
CI - - - - - - - Chronos
CI - - - - - - - LLM4TS
CI - - - - - - - Time-LLM
CI - - - - - - - RevIN
CD CNN-based $\checkmark$ - - - - - Informer
CD CNN-based $\checkmark$ - - - - - Autoformer
CD CNN-based $\checkmark$ - - - - - FEDformer
CD CNN-based $\checkmark$ - - - - - TimesNet
CD MLP-based $\checkmark$ - - - - - TSMixer
CD MLP-based $\checkmark$ - - - - - TTM
CD Transformer-based $\checkmark$ - - - $\checkmark$ - iTransformer
CD Transformer-based $\checkmark$ - - - $\checkmark$ - Crossformer
CD Transformer-based $\checkmark$ $\checkmark$ - - - - VCformer
CD Transformer-based $\checkmark$ $\checkmark$ - - $\checkmark$ - MOIRAI
CD Transformer-based $\checkmark$ - - - $\checkmark$ - UniTS
CD GNN-based - - - $\checkmark$ - - GTS
CD GNN-based $\checkmark$ - - - - $\checkmark$ MSGNet
CD GNN-based - $\checkmark$ - - - - FourierGNN
CD GNN-based - $\checkmark$ - - - - FC-STGNN
CD GNN-based $\checkmark$ - - - $\checkmark$ - TPGNN
CD Linear Channel Mixing (Factorization) $\checkmark$ - - - - - SOFTS
CD Others $\checkmark$ - - - $\checkmark$ - C-LoRA
CP CNN-based $\checkmark$ - - - - - ModernTCN
CP Transformer-based $\checkmark$ - - $\checkmark$ - - DUET
CP Transformer-based $\checkmark$ - - - $\checkmark$ - MCformer
CP Transformer-based $\checkmark$ - - $\checkmark$ $\checkmark$ - DGCformer
CP Transformer-based - - - - $\checkmark$ - CM
CP GNN-based $\checkmark$ - - - - - MTGNN
CP GNN-based $\checkmark$ - $\checkmark$ - - - CrossGNN
CP GNN-based $\checkmark$ - - $\checkmark$ - - WaveForM
CP GNN-based - - - - $\checkmark$ - MTSF-DG
CP GNN-based $\checkmark$ - - $\checkmark$ - - ReMo
CP GNN-based $\checkmark$ - - $\checkmark$ $\checkmark$ $\checkmark$ Ada-MSHyper
CP Others $\checkmark$ $\checkmark$ - - - - LIFT
CP Others $\checkmark$ - - $\checkmark$ - - CCM
CI + CD Parallel CNN (CI) and BiLSTM/Transformer(CD) based Dual Encoder + ETA $\checkmark$ - - - $\checkmark$ $\checkmark$ MSTN
The established field of multivariate time–series forecasting is conventionally organized around three channel interaction strategies—channel–independent (CI), channel–dependent (CD), and channel–projected (CP) approaches [qiu2025comprehensivesurveydeeplearning]—as summarized in Table 6. These strategies specify how a model encodes and exploits inter-variable dependencies across the $C$ channels of a multivariate signal.
CI. CI methods (e.g., DLinear, PatchTST) process each variable independently using lightweight temporal operators such as linear layers or multilayer perceptrons. Their advantage lies in linear channel cost $\mathcal{O}(C)$ and efficient temporal complexity $\mathcal{O}(L)$, though this comes at the expense of discarding cross-channel structure.
CD. CD approaches (e.g., iTransformer) explicitly model inter-variable correlation through dense attention or graph-based mechanisms. While these models offer comprehensive feature coupling, they incur high computational cost—typically $\mathcal{O}(C^{2})$ or $\mathcal{O}(L^{2})$—which limits their scalability and latency characteristics.
CP. CP models (e.g., MCformer, MTGNN) map the channel dimension into a low-rank latent space before applying interaction, thereby avoiding the full $\mathcal{O}(C^{2})$ cost of CD models. Although channel-efficient, these models often impose simplified temporal assumptions, reducing their ability to capture multi-scale dynamics.
Despite the usefulness of this taxonomy, it does not fully resolve the central challenges in long-sequence time–series applications, where models must simultaneously capture short- and long-range temporal dependencies while ensuring low-latency inference. Consequently, there is a need for architectures capable of representing multi-scale temporal structure without incurring prohibitive computational overhead.
Existing SOTA systems illustrate these trade-offs. CI-based PatchTST [nie2023timeseriesworth64] relies on fixed patch-based tokenization; LLM-derived approaches such as LLM4TS [10.1145/3719207] and TIME-LLM [jin2024timellmtimeseriesforecasting] depend on frozen backbone architectures; CD-based TimesNet [wu2023timesnettemporal2dvariationmodeling] employs fixed frequency priors; and CP-based MCformer [10533212] attains efficiency through partial channel blending. These designs incorporate rigid structural priors that limit their ability to capture diverse temporal scales without sacrificing computational efficiency.
To address these limitations, we introduce MSTN, a hybrid architecture grounded in the Efficient Temporal Aggregation (ETA) principle. MSTN explicitly targets the multi-scale nature of long-sequence data while avoiding the $\mathcal{O}(L^{2})$ bottleneck typical of Transformer-based models, enabling $\mathcal{O}(1)$ inference with respect to sequence length. By combining the complementary strengths of CI and CD models with a novel temporal aggregation mechanism, MSTN delivers strong performance across forecasting, imputation, and classification tasks while maintaining the low-latency characteristics required for real-world deployment.
4.2.1 Core Innovation: Dual-Path Design and ETA
Table 7: Comparison of architectural characteristics. This table contrasts the MSTN framework against the three standard Channel Interaction strategies (CI, CD, CP) across critical performance dimensions. MSTN achieves high capacity while neutralizing the $\mathcal{O}(L^{2})$ computational bottleneck.
Dimension CI CD CP (CI+CD) MSTN Rationale
Efficiency High Low Moderate High ETA enables $\mathcal{O}(1)$ refinement after initial encoding.
Robustness High Low Moderate High Dual-encoder with SGF provides complementary feature stability.
Generaliza- bility Low Moderate High High Achieved via a single, aggregated feature vector ($\mathbf{z}_{\text{final}}$) that supports multi-task learning (Forecasting, Imputation, Classification).
Capacity Low High Moderate High Transformer/BiLSTM path captures global dependencies.
Ease of Implementation High Moderate Low Moderate Relies on standard components (CNN, Transformer/BiLSTM, SGF, SE, MHA) rather than complex, specialized graph or recursive structures.
The MSTN framework integrates a dual-path temporal encoding architecture with the ETA mechanism, yielding a principled hybrid approach that reconciles the core trade-offs among CI-, CD-, and CP-based modelling strategies as shown in Table 7. Its design is centered on two architectural innovations that jointly enhance representational fidelity and computational efficiency:
  1. 1.
    Dual-branch temporal encoding. In contrast to recent CP-oriented architectures such as MCformer [10533212], which prioritise efficiency at the expense of temporal expressiveness, MSTN explicitly preserves multi-scale temporal structure. It deploys a parallel encoding strategy in which a global sequence-modeling pathway (Transformer or BiLSTM; CD strategy) captures long-range temporal dependencies, while a lightweight convolutional pathway (CI strategy; $\mathcal{O}(L)$) extracts fine-grained local temporal patterns. This duality directly addresses the multi-scale nature of real-world temporal signals, offering coverage beyond what is typically afforded by efficient CP-based methods.
  2. 2.
    Efficiency through ETA: collapsing $\mathcal{O}(L^{2})$ to $\mathcal{O}(1)$. MSTN applies the ETA mechanism immediately after the high-capacity encoders, collapsing the temporal dimension ($L\to 1$) via learned sequence aggregation. By performing the computationally intensive operations (e.g., $\mathcal{O}(L^{2})$ Transformer self-attention) before this aggregation step, the subsequent refinement layers—including feature fusion, squeeze–excitation recalibration, multi-head attention, and prediction modules—operate with a fixed $\mathcal{O}(1)$ cost with respect to $L$. This explicitly reshapes the model’s complexity profile, enabling MSTN to maintain the representational power of CD architectures while achieving substantially improved inference efficiency.
4.3 Baselines
This section presents the baselines used to assess MSTN on four time series tasks: (1) imputation, (2) long-term forecasting, (3) classification, and (4) generalizability study.
4.3.1 Imputation
Following the TimesNet [wu2023timesnettemporal2dvariationmodeling] protocol, we evaluate performance under the MCAR scenario with random masking ratios of $\{12.5\%,25\%,37.5\%,50\%\}$. Performance is quantified using MSE and MAE. The evaluation includes recent SOTA baselines such as GPT2(3) [zhou2023fitsallpowergeneraltime], TimesNet [wu2023timesnettemporal2dvariationmodeling], PatchTST [nie2023timeseriesworth64], LightTS [zhang2022morefastmultivariatetime], and DLinear [zeng2022transformerseffectivetimeseries].